MODS1 Commissioning Run 4: UTC 2011 January 13

Observer(s): R. Pogge (OSU), M. Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
AO Support: D. Miller (LBTO), from Tucson as required
Telescope Support: J. Little (LBTO)
Instrument Support: T. Shih (LBTO)

Summary

Work continues on the bogie bearings during the day. Skies are clear now, so we wish the engineers well.

Glycol flow was cutoff to the instrument without warning during the morning by work on the telescope. MODS Environmental Monitor logs show the cutoff occurred at 1629UTC (9:29am MST). The CCD controller head electronics boxes (HEBs) thermally protected themselves as they should under such circumstances (depending on the ambient temperature, they will heat up to their auto shutdown temperature in 30-40 minutes after glycol flow stops). The cause was work on a UPS connected to telescope systems that include the instrument cooling system.

Instrument cooling service was restored at about 2025UTC. After waiting for it to stabilize, the CCD controllers were restarted without incident.

The telescope is still not functional - the problem with the bogies gets more complex as they continue to work on it. There will be a review in Tucson tomorrow regarding the next steps.

Preparations

New MODS mmcServer code loaded to implement new mid-level commands for switching efficiently between calibration and observing modes. Will test it early this evening.

We will coordinate with the AO team on the mountain regarding testing this evening since observing is not possible.

Details

All times are UTC unless otherwise noted.

-- RichardPogge - 12 Jan 2011
Topic revision: r3 - 13 Jan 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback